# A Monte Carlo Tree Search approach to QAOA: finding a needle in the haystack

### v2

@@ III.1 The 3-Satisfiability problem
[…]
@@
[…]
In this work, we work with a set of 3-SAT instances with $\alpha=3$, each with a unique bit-string satisfying Eq. (4). Despite this density of clauses being lower than the threshold for the complexity phase transition in the thermodynamic limit [58], having a non-degenerate ground state makes the optimization hardness comparable to random problem instances in the critical regions [63, 36]. From a quantum annealing perspective, this is reflected in the quantum Hamiltonian having an energy gap that decreases exponentially with the qubit number, making annealing unfeasible for large systems. 3-SAT is known to be a formidable problem for QAOA, requiring large circuit depths($P\gtrsim 45$ for 0.95 fidelity on 6 qubits [64]) to reach the ground state with high fidelity. However, as long as the projection over the ground state is large enough, an approximate optimization is sufficient to sample it with a finite probability. As deep variational circuits are unfeasible for current hardware, we will limit the number of layers to $P\leq 10$. As a mixer Hamiltonian for QAOA we use a transverse field
@@
Figure 2: Comparison of the performance of different versions of MCTS on a $n=7$ example 3-SAT instance. Energies of the final states are shown as a function of circuit depth for vanilla MCTS (blue), MCTS with single-player modifications (orange), iterative SSR-MCTS (green), and iterative SSR-MCTS with single-player modifications (green). The number of MCTS cycles increases linearly with $P$.
@@ III.2 Unrestricted MCTS VS SSR-MCTS
Figure 3: Distribution of all leaf nodes in the tree representing all possible combinations of parameters MCTS can choose for a circuit depth of $P=2$ and a branching factor $b=30$, in an example 3-SAT instance. (a) Leaf nodes from the full parameter space (up to inversion symmetry), (b) leaf nodes within the restricted parameter space defined according to the $P=1$ optimal angles. Each leaf node is represented according to its difference in energy and tree distance (defined in App. B.1) from the optimal node $(F_{2}(\bm{\theta})-F_{2}(\bm{\theta}^{*}),d(\bm{\theta},\bm{\theta}^{*}))$, and the color of each leaf node represents the sum of the corresponding QAOA parameters (run time). The nodes corresponding to the best possible sets of parameters are the ones on (0,0), the bottom-left corners of the figures. For each of the two cases, we also show the five leaf nodes selected by both standard MCTS (blue crosses) and MCTS-SP (orange diamonds) after five runs, each with different numbers of MCTS cycles per turn. The symbols in the squares correspond to the results with the fewest amount of cycles, and the lines connect the leaf nodes in order of increasing cycle numbers in the set $\{200,600,1000,1800,3500\}$.
The dramatic effect of SSR already at $P=2$ is due to its impact on the leaf-node structure. To appreciate it, in Fig. 3 we compare the energy distribution among the leaf nodes for a depth $P=2$ when MCTS has the unrestricted $[0,2\pi)^{2P}$ space to search in (a), with SSR generated from the $P=1$ optima in (b). Each leaf node can be identified by a tuple composed by its energy difference to the optimal choice $F_{P}(\bm{\theta})-F_{P}(\bm{\theta}^{*})$, the corresponding tree distance $d(\bm{\theta},\bm{\theta}^{*})$ (a measure of the deviation from the path of the optimal leaf, defined in App. B.1), and the effective run time $\tau=\sum_{i}(|\gamma_{i}|+|\beta_{i}|)$. Figure 3 shows the relation between these quantities.
@@
If we look at SSR in panel (b), instead, a large energy difference also corresponds to a node far from the optimum. This means that good paths down the tree are more concentrated in particular branches and can be reached via similar parameter choices in the early turns of the game, making MCTS able to discover and learn strategies to navigate the cost landscape adeptly. The energy interval spanned by the possible QAOA schedules is also reduced, notice the difference scales on the $x$-axis. SSR gets rid of all terminal nodes corresponding to excessively large run times and highly excited states, thus focusing on the parameter region where fast-running and low-energy minima reside. More details on the leaf node structure are given in Appendix B.
The difficulties in vanilla MCTS persist if one tries to improve its performance by increasing the number of cycles. An example is provided by the set of blue crosses in both panels of Fig. 3. Starting from 200 MCTS cycles per turn (boxed cross) we increase this number to $\{600,1000,1800,3500\}$. The links connect MCTS results with a growing number of cycles. When MCTS has to search through the whole parameter landscape, panel (a), it typically gets stuck on leaf nodes far away from the optimal one. On the contrary, it consistently finds well-scoring nodes near the optimal one for the SSR case.
@@
Now that we had a glimpse into the importance of search-space restriction for transforming the QAOA decision tree into one that MCTS can efficiently find optimal strategies in, let us delve into more details of the iterative SSR-MCTS protocol we use. In each run, the algorithm performs 1000 initial cycles plus 800 more per turn, excluding the final one, while selecting the parameters. The last parameter to choose does not require further stochastic rollouts since the best move can be selected straightforwardly. Hence, even though the search space scales exponentially with $P$, the number evaluations of $F_{P}(\bm{\gamma},\bm{\beta})$ only increases linearly with growing circuit depth, $n_{\rm fev}=1000+800(2P-1)$, as two new turns are added at each subsequent $P$. See appendix D for an analysis of the performance dependence on the number of training cycles. We set $C=\sqrt{2}$ in the selection policy and $\nu=1/2$ for the exponential mapping between energy and reward discussed in II.2. At each depth $P$, the range of the parameters is bounded according to the parameters proposed for the previous depth, and a branching factor of $b=30$ is used to build the tree within the restricted space. The boundaries are softened with the values $\bm{\delta}=\{0,0,0.1,0.05,0.04,0.03,0.02,0.01,0.01\}$, where $\delta_{P}$ indicates how much the search is allowed to search outside of the regular regions at depth $P$, according to (II.3). We set $\delta_{1}=0$ since we are already considering the entire space at the shallowest depth, and $\delta_{2}=0$ because we expect the algorithm to find the $P=1$ optimal parameters in every case. We set a nonlinear decrease of $\delta_{P}$ with the circuit depth when $P\geq 2$.
Figure 4: Average energies obtained with the iterative SSR-MCTS algorithm for different problem sizes for 3-SAT as a function of the circuit depth. The energies for each $n$ are averaged over 15 problem instances, each with a clause to variable ratio $\alpha=3$ and one unique satisfying ground state. The algorithm performs $1000+800(2P-1)$ cycles at depth $P$.
@@
The effect of SSR on the QAOA schedule can be appreciated from Fig. 5, where we show an example set of parameters chosen by iterative SSR-MCTS for different depths on a 3-SAT instance with $n=7$. Each dataset is a sequence $(\beta_{i},\gamma_{i})$ with the first pair corresponding to the point closer to the lower-right corner of the plot. As $P$ increases, the optimal parameters clearly approach a regular behavior, linked to optimal digitized-quantum-annealing schedules and known to be associated with optimal minima strategies robust with respect to small parameter and size variations [9, 19, 28]. While iterative SSR is the analog of other iterative optimization methods of QAOA parameters [9, 19] in the context of MCTS, it is remarkable that it completely changes standard MCTS performance already at $P=2$. While the regularity at large circuit depth is expected, its appearance is noteworthy when the search space is only moderately restricted.
Figure 5: Example QAOA angles chosen by iterative SSR-MCTS to approximate the ground state of an example $n=9$ 3-SAT instance at different circuit depths. The parameters from $P=1$ to $P=5$ are shown on the left, and the $P=9$ to $P=10$ ones on the right. The visualization of the parameters is similar to that of Fig. 1, with the parameter indices increasing from bottom-right corners to top-left ones.
@@ III.4 Noisy rewards
We model the measurement uncertainty by introducing a random Gaussian variable to the rewards during the rollout phase of each MCTS cycle. Consequently, the algorithm must select variational parameters based on noisy data accumulated in the statistics of the decision tree. Specifically, we use random variables $\mathcal{N}(0,n_{s})$ with zero mean and a standard deviation determined by the noise level $n_{s}$. This addition modifies the reward function as $F_{P}(\bm{\theta})\rightarrow F_{P}(\bm{\theta})+\mathcal{N}(0,n_{s})$. For evaluation purposes, the final assessment does not include this random noise. MCTS generally handles noisy rewards well, though its effectiveness can vary based on factors like noise intensity, problem structure, and algorithm implementation [41]. The relevant energy scale for comparing noise strength is the typical energy gap $\Delta$ between eigenvalues. As long as $n_{s}\ll\Delta$, MCTS should effectively distinguish between leaf nodes leading to states close to different eigenstates, as the node scores are already subject to uncertainty due to stochastic sampling. However, if the noise strength approaches the gap $n_{s}\sim\Delta$, it can significantly affect the estimation of optimal actions, making it difficult to differentiate between strategies. In our 3-SAT formulation (see Eq. (5)), the energy units correspond to the number of violated clauses, with the minimum relevant energy gap being $\Delta=1$.
To test the protocol’s robustness against noise, we focus on the hardest 3-SAT instance (smallest energy gap in a QA setting) with $n=7$ and repeat 15 SSR-MCTS runs, without the single-player modifications (see App. C for a comparison with other variants and COBYLA [69]), for 3 different jumps on $P$ between 1 and 8, and $n_{s}\in[0,1.5]$. Remarkably, MCTS has no issues finding the optimal parameters as it is able to average out the noise over different cycles even for large noise strengths $n_{s}\gtrsim\Delta$, as we show in Fig. 6. As the noise strength is increased, the average performance gets mildly worse (full circles with solid lines) but it is often able to find very good parameters in one of the runs. Indeed, the best results out of the 15 repetitions (crosses with dashed lines) drift away even more slowly than the average from the noise-free QAOA energy, as $n_{s}$ increases. This remarkable robustness makes iterative SSR-MCTS well-suited for applications in noisy quantum devices, as it does not rely on gradient estimations and the uncertainty on the reward estimate can be easily taken into account when designing how the leaf nodes are sampled.
@@ IV Results - Unweighted MaxCut
[…]
@@
such that $r=1$ if the optimal cut is found. In Fig. 7 we show the results of iterative SSR-MCTS for 19 MaxCut problem instances on $n=10$ unweighted 3-regular graphs. We plot both the energy for each individual instance (blue crosses) and the average approximation ratio over all instances (orange circles). The latter is compared to the performance guarantee of the Goemans-Williamson (GW) algorithm [70, 71], which is a standard classical reference for MaxCut problems. The convergence to the ground state is smooth with increasing values of $P$, similar to what we observed for 3-SAT problems, and the performance is qualitatively similar to the results of Zhou et al. [19]. At $P=4$, iterative SSR-MCTS optimization of QAOA schedules beats, on average, the GW algorithm, confirming again that the performance is comparable with standard QAOA results. Notice though, that for $E\to 0$ (or $r\to 1$), the discretization of the parameter space required for MCTS might prevent the algorithm from finding the exact ground state ($r=1$) even if it is reachable by the variational QAOA ansatz. Improved search-space restriction schemes for the edge parameters, such as searching a limited distance from the previous optima as suggested in Sec. II.3 or making the discretization non-uniform (finer closer to the previous optima), could be a first step in fully optimizing the performance of iterative SSR-MCTS for deep circuits. In any case, this does not pose a limitation in experiments, as the hardware noise would likely be larger, and sampling the ground state with high probability does not require perfect fidelity with the final variational state.

### v3

@@ III.1 The 3-Satisfiability problem
As a first benchmark for the MCTS-guided quantum optimization, we consider instances of boolean 3-Satisfiability (3-SAT) problems. 3-SAT is one of the most well-known $\mathsf{NP}$-complete problems [62] and a standard benchmark for quantum optimization algorithms [63, 39]. Given a set of $n$ classical binary variables $x_{i}\in\{0,1\}$, it consists in finding whether a particular assignment $\bm{x}^{*}$ exists such that a boolean expression consisting of a logical conjunction of $m$ clauses
@@
The expectation value of the cost Hamiltonian Eq. (5) gives the average number of unsatisfied clauses in the state under consideration. Finding the ground state of said Hamiltonian is thus equivalent to solving the $\mathsf{NP}$-hard optimization version of the corresponding 3-SAT problem, also known as MAX-3-SAT [62], where the (harder) task is finding the bitstring that violates the minimum number of clauses. Clearly, this also solves 3-SAT, as the problem is satisfiable if and only if the ground state energy is zero. The phase transition of MAX-3-SAT happens for similar clause densities to $\alpha_{c}$, although it affects the hardness of the problem differently  [^5] The hardness of MAX-3-SAT also depends on the clause density $\alpha$, and one expects a phase transition to occur for some $\alpha\leq\alpha_{c}$, as MAX-3-SAT subsumes the complexity of 3-SAT. As $\alpha$ increases past $\alpha_{c}$ and the problem instances become more and more over-constrained, the hardness of MAX-3-SAT, unlike that of 3-SAT, never decreases. Such high clause densities make it easy to find that problem instances are unsatisfiable, but very hard to find their optimal bitstrings. We have gone from an easy-hard-easy type of transition for 3-SAT to an easy-hard one for MAX-3-SAT [65]. Even though we are technically solving MAX-3-SAT, in the following we will refer to the problem instances as “3-SAT instances”, as it is customary for the sake of convenience.
In this work, we work with a set of 3-SAT instances with $\alpha=3$, each with a unique bit-string satisfying Eq. (4). Despite this density of clauses being lower than the threshold for the complexity phase transition in the thermodynamic limit [64], having a non-degenerate ground state makes the optimization hardness comparable to random problem instances in the critical regions [69, 39]. From a quantum annealing perspective, this is reflected in the quantum Hamiltonian having an energy gap that decreases exponentially with the qubit number, making annealing unfeasible for large systems. 3-SAT is known to be a formidable problem for QAOA, requiring large circuit depths to reach the ground state with high fidelity ($P\gtrsim 45$ for 0.95 fidelity on 6 qubits [70]). However, as long as the projection over the ground state is large enough, an approximate optimization is sufficient to sample it with a finite probability. As deep variational circuits are unfeasible for current hardware, we will limit the number of layers to $P\leq 10$. As a mixer Hamiltonian for QAOA we use a transverse field
@@
Figure 2: Comparison of the performance of different versions of MCTS on a $n=7$ example 3-SAT instance. Energies of the final states are shown as a function of circuit depth for vanilla MCTS (blue), MCTS with single-player modifications (orange), iterative SSR-MCTS (green), and iterative SSR-MCTS with single-player modifications (red). The number of MCTS cycles increases linearly with $P$.
@@ III.2 Unrestricted MCTS VS SSR-MCTS
Figure 3: Distribution of all leaf nodes in the tree representing all possible combinations of parameters MCTS can choose for a circuit depth of $P=2$ and a branching factor $b=30$, in an example 3-SAT instance. (a) Leaf nodes from the full parameter space (up to inversion symmetry), (b) leaf nodes within the restricted parameter space defined according to the $P=1$ optimal angles. Each leaf node is represented according to its difference in energy and tree distance (defined in B.1) from the optimal node $(F_{2}(\bm{\theta})-F_{2}(\bm{\theta}^{*}),d(\bm{\theta},\bm{\theta}^{*}))$, and the color of each leaf node represents the sum of the corresponding QAOA parameters (run time). The nodes corresponding to the best possible sets of parameters are the ones on (0,0), the bottom-left corners of the figures. For each of the two cases, we also show the five leaf nodes selected by both standard MCTS (blue crosses) and MCTS-SP (orange diamonds) after five runs, each with different numbers of MCTS cycles per turn. The symbols in the squares correspond to the results with the fewest number of cycles, and the lines connect the leaf nodes in order of increasing cycle numbers in the set $\{200,600,1000,1800,3500\}$.
The dramatic effect of SSR already at $P=2$ is due to its impact on the leaf-node structure. To appreciate it, in Fig. 3 we compare the energy distribution among the leaf nodes for a depth $P=2$ when MCTS has the unrestricted $[0,2\pi)^{2P}$ space to search in (a), with SSR generated from the $P=1$ optima in (b). Each leaf node can be identified by a tuple composed by its energy difference to the optimal choice $F_{P}(\bm{\theta})-F_{P}(\bm{\theta}^{*})$, the corresponding tree distance $d(\bm{\theta},\bm{\theta}^{*})$ (a measure of the deviation from the path of the optimal leaf, defined in B.1), and the effective run time $\tau=\sum_{i}(|\gamma_{i}|+|\beta_{i}|)$. Figure 3 shows the relation between these quantities.
@@
If we look at SSR in panel (b), instead, a large energy difference also corresponds to a node far from the optimum. This means that good paths down the tree are more concentrated in particular branches and can be reached via similar parameter choices in the early turns of the game, making MCTS able to discover and learn strategies to navigate the cost landscape adeptly. The energy interval spanned by the possible QAOA schedules is also reduced, notice the difference between the scales on the $x$-axis. SSR gets rid of all terminal nodes corresponding to excessively large run times and highly excited states, thus focusing on the parameter region where fast-running and low-energy minima reside. More details on the leaf node structure are given in Appendix B.
The difficulties of vanilla MCTS persist if one tries to improve its performance by increasing the number of cycles. An example is provided by the set of blue crosses in both panels of Fig. 3. Starting from 200 MCTS cycles per turn (boxed cross) we increase this number to $\{600,1000,1800,3500\}$. The links connect MCTS results with a growing number of cycles. When MCTS has to search through the whole parameter landscape, panel (a), it typically gets stuck on leaf nodes far away from the optimal one. On the contrary, it consistently finds well-scoring nodes near the optimal one for the SSR case.
@@
Now that we had a glimpse into the importance of search-space restriction for transforming the QAOA decision tree into one that MCTS can efficiently find optimal strategies in, let us delve into more details of the iterative SSR-MCTS protocol we use. In each run, the algorithm performs 1000 initial cycles plus 800 more per turn, excluding the final one, while selecting the parameters. The last parameter to choose does not require further stochastic rollouts since the best move can be selected straightforwardly. Hence, even though the search space scales exponentially with $P$, the number evaluations of $F_{P}(\bm{\gamma},\bm{\beta})$ only increases linearly with growing circuit depth, $n_{\rm fev}=1000+800(2P-1)$, as two new turns are added at each subsequent $P$. See Appendix D for an analysis of the performance dependence on the number of training cycles. We set $C=\sqrt{2}$ in the selection policy and $\nu=1/2$ for the exponential mapping between energy and reward discussed in II.2. At each depth $P$, the range of the parameters is bounded according to the parameters proposed for the previous depth, and a branching factor of $b=30$ is used to build the tree within the restricted space. The boundaries are softened with the values $\bm{\delta}=\{0,0,0.1,0.05,0.04,0.03,0.02,0.01,0.01\}$, where $\delta_{P}$ indicates how much the search is allowed to search outside of the regular regions at depth $P$, according to (II.3). We fix $\delta_{1}=0$ since we are already considering the entire space at the shallowest depth, and $\delta_{2}=0$ because we expect the algorithm to find the $P=1$ optimal parameters in every case. We set a nonlinear decrease of $\delta_{P}$ with the circuit depth when $P\geq 2$.
Figure 4: (a) Average energies obtained with the iterative SSR-MCTS algorithm for different problem sizes for 3-SAT as a function of the circuit depth. The energies for each $n$ are averaged over 15 problem instances, each with a clause to variable ratio $\alpha=3$ and one unique satisfying ground state. The algorithm performs $1000+800(2P-1)$ cycles at depth $P$. (b) Example QAOA angles chosen by iterative SSR-MCTS to approximate the ground state of an example $n=9$ 3-SAT instance at different circuit depths. The parameters from $P=1$ to $P=5$ are shown on the top panel, and the $P=9$ to $P=10$ ones on the bottom panel. The visualization of the parameters is similar to that of Fig. 1, with the parameter indices increasing from bottom-right corners to top-left ones.
@@
The effect of SSR on the QAOA schedule can be appreciated from Fig. 4(b), where we show an example set of parameters chosen by iterative SSR-MCTS for different depths on a 3-SAT instance with $n=7$. Each dataset is a sequence $(\beta_{i},\gamma_{i})$ with the first pair corresponding to the point closer to the lower-right corner of the plot. As $P$ increases, the optimal parameters clearly approach a regular behavior, linked to optimal digitized-quantum-annealing schedules and known to be associated with optimal minima strategies robust with respect to small parameter and size variations [11, 21, 30]. While iterative SSR is the analog of other iterative optimization methods of QAOA parameters [11, 21] in the context of MCTS, it is remarkable that it completely changes standard MCTS performance already at $P=2$. While the regularity at large circuit depth is expected, its appearance is noteworthy when the search space is only moderately restricted.
@@ III.4 Noisy rewards
We model the measurement uncertainty by introducing a random Gaussian variable to the rewards during the rollout phase of each MCTS cycle. Consequently, the algorithm must select variational parameters based on noisy data accumulated in the statistics of the decision tree. Specifically, we use random variables $\mathcal{N}(0,n_{s})$ with zero mean and a standard deviation determined by the noise level $n_{s}$. This addition modifies the reward function as $F_{P}(\bm{\theta})\rightarrow F_{P}(\bm{\theta})+\mathcal{N}(0,n_{s})$. For evaluation purposes, the final assessment does not include this random noise. MCTS generally handles noisy rewards well though its effectiveness can vary based on factors like noise intensity, problem structure, and algorithm implementation [44]. The relevant energy scale for comparing noise strength is the typical energy gap $\Delta$ between eigenvalues. As long as $n_{s}\ll\Delta$, MCTS should effectively distinguish between leaf nodes leading to states close to different eigenstates, as the node scores are already subjected to uncertainty due to stochastic sampling. However, if the noise strength approaches the gap $n_{s}\sim\Delta$, it can, in principle, significantly affect the estimation of optimal actions, making it difficult to differentiate between strategies. In our 3-SAT formulation (see Eq. (5)), the energy units correspond to the number of violated clauses, with the minimum relevant energy gap being $\Delta=1$.
To test the protocol’s robustness against noise, we focus on the hardest 3-SAT instance (smallest energy gap in a QA setting) with $n=7$ and repeat 15 SSR-MCTS runs, without the single-player modifications (see Appendix C for a comparison with other variants and COBYLA [75]), for 3 different jumps on $P$ between 1 and 8, and $n_{s}\in[0,1.5]$. Remarkably, MCTS has no issues finding the optimal parameters as it is able to average out the noise over different cycles even for large noise strengths $n_{s}\gtrsim\Delta$, as we show in Fig. 5. As the noise strength is increased, the average performance gets mildly worse (full circles with solid lines) but it is often able to find very good parameters in one of the runs. Indeed, the best results out of the 15 repetitions (crosses with dashed lines) drift away even more slowly than the average from the noise-free QAOA energy, as $n_{s}$ increases. This remarkable robustness makes iterative SSR-MCTS well-suited for applications in noisy quantum devices, as it does not rely on gradient estimations and the uncertainty on the reward estimate can be easily taken into account when designing how the leaf nodes are sampled.
@@ IV Results - Unweighted MaxCut
To confirm that our results are not unique to 3-SAT, we repeat the performance analysis on unweighted MaxCut instances. MaxCut is another prototypical $\mathsf{NP}$-hard optimization problem, extensively studied in QAOA contexts [18, 21, 78, 79, 80]. It consists in classifying the vertices $V$ of a graph $G(V,E)$ into two sets such that the number of edges in $E$ joining vertices between the different sets is maximal. The problem Hamiltonian for MaxCut can be expressed as an antiferromagnetic Ising model on $G$,
@@
such that $r=1$ if the optimal cut is found. In Fig. 6 we show the results of iterative SSR-MCTS for 19 MaxCut problem instances on $n=10$ unweighted 3-regular graphs. We plot both the energy for each individual instance (blue crosses) and the average approximation ratio over all instances (orange circles). The latter is compared to the performance guarantee of the Goemans-Williamson (GW) algorithm [76, 77], which is a standard classical reference for MaxCut problems. The convergence to the ground state is smooth with increasing values of $P$, similar to what we observed for 3-SAT problems.
The accuracy of iterative SSR-MCTS is qualitatively similar to the results obtained with warm-start gradient descent approaches [21]. At $P=4$, both optimization strategies of QAOA schedules beat, on average, the GW algorithm. At larger values of $P$, when $E\to 0$, the discretization of the parameter space marginally limits the performance of our method, preventing the algorithm from reaching the exact ground state ($r=1$) even if reachable by the variational QAOA ansatz. For instance, at $n=10$ and $P=10$, we find $1-r\simeq 0.01$, one order of magnitude larger than Ref. [21]. A larger branching factor in the decision tree would result in a finer resolution of the parameters, which mitigates this limitation at very small residual energies, at the cost of making the tree exploration more resource-expensive. Furthermore, improved search-space restriction schemes for the edge parameters, such as searching a limited distance from the previous optima as suggested in Sec. II.3 or making the discretization non-uniform (finer closer to the previous optima), could be another step in fully optimizing the performance of iterative SSR-MCTS for deep circuits. In general, however, we can not expect MCTS to outperform gradient-based algorithms working with continuous parameters in ideal simulations where the cost-function is noiseless and the energy landscape has only a few smooth minima, such as the regular MaxCut instances tested here. In experiments, instead, it can be advantageous because of its resilience against unavoidable noise. Furthermore, with single-shot measurements, sampling the ground state with high probability does not require perfect fidelity with the final variational state.